OpenStudy (anonymous):

A 6-sided die is rolled three times. What is the probability that all three rolls resulted in different numbers? 0.500 0.556 0.578 0.596 0.606

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (zehanz):

The first time it may be any number. The second time, one of the numbers shouldn't come up. Probability 5/6. Third time: only 4 possibilities left, so 4/6. Total probability: \(1 \cdot \frac{5}{6}\cdot \frac{4}{6}=...\)

OpenStudy (anonymous):

5/9?

OpenStudy (anonymous):

b

OpenStudy (zehanz):

First multiply, then simplify: 20/36=10/18=5/9

OpenStudy (zehanz):

So, yes it's b.